# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ D.2 Derivation of $\boldsymbol{p_{\mu_{c}}\left({x}\right)}$ for some emblematic cases
$\begin{split}&\left\langle{\rho^{(1;1)}}\right\rangle=\left\langle{\rho^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}x\,p_{g^{(1)}_{j}}^{(\chi)}\left({x}\right)dx=\int_{\mathbb{R}}x\left(\frac{1}{2}\delta(x)+\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\right)dx=\\ &\int_{0}^{+\infty}x\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)dx=\sigma_{w}^{2}\int_{0}^{+\infty}\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)d\frac{x^{2}}{2\sigma_{w}^{2}}=\frac{\sigma_{w}}{\sqrt{2\pi}}.\end{split}$ (81)
The first notable difference in the case of ReLU activation is that $\left\langle{\rho^{(1;1)}}\right\rangle\neq 0$, thus indicating the emergence of IGB. Specifically, from Eq. (74), it follows:
$p_{\mu_{c}}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{2}\frac{\sigma_{w}^{2}}{2\pi}\right)=\mathcal{N}\left(x;0,\frac{2}{\pi}\right),$ (82)
@@
In this scenario, unlike the linear activation function case discussed at the end of App. D.2, we do not need to consider the finite size effects of the dataset. While for a linear activation function $\lim_{D\rightarrow\infty}\left\langle{\rho^{(1;1)}}\right\rangle=0$, for ReLU $\lim_{D\rightarrow\infty}\left\langle{\rho^{(1;1)}}\right\rangle\neq 0$ (as shown in (81)). Therefore, the finite size corrections of $\mathcal{O}\left(\frac{1}{D}\right)$ are negligible when $D\left\langle{\rho^{(1;1)}}\right\rangle\gg 1$.
Figure D.1: Comparison of the distribution $p_{\mu_{c}}\left({x}\right)$ between the linear activation function (upper plot) and ReLU (bottom plot) across different dataset sizes ($D$). Both plots also include the theoretical distributions derived in App. D.2. Observe the distinct behaviors in the two cases: for the linear activation function, the distribution narrows as $D$ increases, while for the ReLU case, it remains stable. This stability in the ReLU case aligns with expectations from Eq. (81) and Eq. (77). These simulations employed the GB dataset and the SHLP model. For more details, refer to App. I.4.
In Fig. D.1, the distributions $p_{\mu_{c}}\left({x}\right)$ along with their respective theoretical curves, i.e., Eq. (78) and Eq. (82), are compared for the two types of activation functions. Consistent with Eq. (78) and Eq. (82), we observe that in the linear case, the distribution narrows around 0 as $D$ increases, whereas in the ReLU case, it remains stable.
@@
With max pooling, we group the set $\{g^{(1)}_{k}\}$ into subgroups and select the maximum value from each, thereby reducing the dimensionality of the layer. Incorporating this into the analysis in App. D.2, we have:
$\rho^{(1;m)}_{l}=\rho\left(\left\{g\left(h^{(1)}_{j}\right)\right\}_{j\in S_{l}^{m}}\right)=\max_{j\in S_{l}^{m}}\{\max\{0,g^{(1)}_{j}\}\}$, (83)
where $S_{l}^{m}$ indicates the $l$-th subgroup of $m$ nodes.
To proceed with the computation from the previous section, we need to derive $p_{\rho^{(1;m)}_{l}}^{(\chi)}\left({x}\right)$. Generally, if $Y\equiv\max\{X_{1},\dots,X_{m}\}$ with $\{X_{i}\}_{i=1}^{m}$ being i.i.d., we have:
@@
$p_{Y}\left({y}\right)=m\,p_{X}\left({y}\right)F_{X}(y)^{m-1}.$ (85)
For our case, we substitute $p_{X}\left({x}\right)$ with the distribution used in App. D.2, i.e.,
$p_{X}\left({x}\right)\equiv p_{g^{(1)}_{j}}^{(\chi)}\left({x}\right)=\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right).$ (86)
@@
$\begin{split}&F_{X}(y)\equiv\mathbb{P}\left(g^{(1)}_{j}\leq y\mid\mathcal{W}\right)=\int_{-\infty}^{y}\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)dx=\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{y}{\sqrt{2\sigma_{w}^{2}}}\right)\Theta(y).\end{split}$ (87)
Applying the Fisher–Tippett–Gnedenko theorem, we can find an asymptotic result for $p_{\rho^{(1;m)}_{l}}^{(\chi)}\left({x}\right)$ as $m\rightarrow\infty$. For general $m$ values, combining all components yields:
$\begin{split}&\left\langle{\rho^{(1;m)}}\right\rangle=\left\langle{\rho^{(1;m)}_{l}}\right\rangle=\int_{\mathbb{R}}x\,p_{\rho^{(1;m)}_{l}}^{(\chi)}\left({x}\right)dx=\int_{0}^{\infty}m\,x\,\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{x}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{m-1}dx.\end{split}$ (88)
This can be approximated by an asymptotic expansion or evaluated numerically. In the case of linear activation with max pooling, a simple Gaussian distribution is substituted for $p_{X}\left({x}\right)$.
@@
$c_{m}(\sigma_{w})\equiv\int_{0}^{\infty}m\,x\,\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{x}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{m-1}dx.$ (89)
@@
$p_{\mu_{c}}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{2}c_{m}(\sigma_{w})^{2}\right).$ (90)
D.3 Derivation of $\boldsymbol{p_{O^{(c)}}^{(\chi)}\left({x}\right)}$ for some emblematic cases
As discussed, $\mu_{c}$ is a r.v. that varies with the set of weights $\mathcal{W}$. In App. D.2, we derived its distribution. In this section, we focus on the distribution of the output, given a specific configuration of the weight set, $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ , for the same cases analyzed in App. D.2.
@@ D.3 Derivation of $\boldsymbol{p_{O^{(c)}}^{(\chi)}\left({x}\right)}$ for some emblematic cases
    $\lim_{N_{1}\rightarrow\infty}p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{4}\right).$ (91)
@@
    $\left\langle{g^{(1)}_{m}}\right\rangle$ $=\frac{\sigma_{w}}{\sqrt{2\pi}},$ (92)
    $\begin{split}\textrm{Var}_{\chi}\left({g^{(1)}_{m}}\right)&=\int_{\mathbb{R}^{+}}x^{2}p_{g^{(1)}_{m}}^{(\chi)}\left({x}\right)dx-\left\langle{g^{(1)}_{m}}\right\rangle^{2}=\\ &=\frac{\sigma_{w}^{2}(\pi-1)}{2\pi},\end{split}$ (93)
    where the symmetry of $p_{h^{(1)}_{m}}^{(\chi)}\left({x}\right)$ was used. Substituting into Eq. (72) and using Eq. (93), we obtain
@@
  • ReLU + max pool: This case is conceptually similar to the ReLU case; the main difference is that we do not have an analytical expression for $\left\langle{\rho^{(1;m)}_{i}}\right\rangle$ and $\textrm{Var}_{\chi}\left({\rho^{(1;m)}_{i}}\right)$. We can numerically compute these cumulants for a fixed $m$, with their expressions given by
    $\begin{split}&\left\langle{\rho^{(1;m)}}\right\rangle=\int_{0}^{\infty}mx\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{x}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{m-1}dx,\end{split}$ (95)
    $\begin{split}&\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)=\left(\int_{0}^{\infty}mx^{2}\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{x}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{m-1}dx\right)-\left\langle{\rho^{(1;m)}}\right\rangle^{2}.\end{split}$ (96)
@@
    $\begin{split}&\lim_{N_{1}\rightarrow\infty}p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\mu_{c},\sigma_{w}^{2}\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)\right),\end{split}$ (97)
@@ D.4 Non identically distributed classes
$\textrm{Var}_{\chi}(O^{(c^{\prime}.c)})=\frac{\sigma_{w}^{4}\sigma_{\xi^{(c)}}^{2}(\pi-1)}{2\pi}$
@@ Appendix E Single hidden layer perceptron
In this section, we utilize the results derived in App. D to deduce our primary goal, $p_{f_{0}}\left({x}\right)$. We specifically focus on a single hidden layer perceptron; in App. G.3 we will extend the results to deep architectures.
E.1 Derivation of $\boldsymbol{p_{f_{0}}\left({x}\right)}$
Consider a binary classification problem using a single hidden layer perceptron. We will use the notation outlined in Fig. A.1. The nodes of the hidden layer, after processing through the activation function and the pooling layer, are connected to two output nodes $\left\{O^{(c)}\right\}_{c=0}^{1}$ via two random vectors $\left\{w^{(1)}_{c\;\boldsymbol{\cdot}}\right\}_{c=0}^{1}$ . Each output node $O^{(c)}\left(\boldsymbol{\xi};\mathcal{W}^{0},w^{(1)}_{c\;\boldsymbol{\cdot}}\right)$ follows the distributions derived in App. D.3.
Note that, given a set $\mathcal{W}$, different output nodes depend on distinct subsets of $\mathcal{W}$, as the final matrix of weights is not shared among all nodes. The notation $O^{(c)}\left(\boldsymbol{\xi};\mathcal{W}^{0},w^{(1)}_{c\;\boldsymbol{\cdot}}\right)$ highlights this aspect, underscoring why the output nodes have different distributions. However, for brevity, we will denote a general dependency on the entire set $\mathcal{W}$, i.e., we will use the notation $O^{(c)}\left(\boldsymbol{\xi};\mathcal{W}\right)$ .
To derive the distribution $p_{f_{0}}\left({x}\right)$ , we proceed as follows:
In a single experiment, the set $\left\{w^{(1)}_{c\;\boldsymbol{\cdot}}\right\}_{c=0}^{1}$ is fixed, allowing us to compute $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ . Recalling that $f_{0}$ represents the fraction of the dataset classified as belonging to class 0, according to the Law of Large Numbers, this fraction will converge to the probability
$\lim_{D\rightarrow\infty}f_{0}\left(\mathcal{W}\right)=\mathbb{P}\left(O^{(0)}>O^{(1)}\mid\mathcal{W}\right)=\mathbb{P}\left(O^{(0)}-O^{(1)}>0\mid\mathcal{W}\right)=\int_{0}^{\infty}p_{\Delta_{O}}^{(\chi)}\left({z}\right)\;dz,$ (99)
where $f_{0}\left(\mathcal{W}\right)$ indicates the fraction $f_{0}$ obtained for the specific fixed configuration $\mathcal{W}$.
$f_{0}$ is a quantity clearly dependent on the set of random variables $\mathcal{W}$. Consequently, it is a random variable itself, varying with the network’s set of weights. Defining $\Delta_{O}=O^{(0)}-O^{(1)}$ , we can express:
$p_{f_{0}}\left({x}\right)=\int_{\Omega}p_{\mathcal{W}}\left({\tilde{\mathcal{W}}}\right)\delta\left(\mathbb{P}\left(\Delta_{O}>0\mid\tilde{\mathcal{W}}\right)-f_{0}\right)d\tilde{\mathcal{W}}$ (100)

### v5

@@ D.2 Derivation of $\boldsymbol{f_{\mu_{c}}\left({m}\right)}$ for some emblematic cases
$\begin{split}&\left\langle{U^{(1;1)}}\right\rangle=\left\langle{U^{(1;1)}_{j}}\right\rangle=\int_{\mathbb{R}}h{}{}\,f_{H^{(1)}_{j}}^{(\chi)}\left({h{}{}}\right)dh{}{}=\int_{\mathbb{R}}h{}{}\left(\frac{1}{2}\delta(h{}{})+\mathcal{N}\left(h{}{};0,\sigma_{w}^{2}\right)\right)dh{}{}=\\ &\int_{0}^{+\infty}h{}{}\mathcal{N}\left(h{}{};0,\sigma_{w}^{2}\right)dh{}{}=\sigma_{w}^{2}\int_{0}^{+\infty}\mathcal{N}\left(h{}{};0,\sigma_{w}^{2}\right)d\frac{h{}{}^{2}}{2\sigma_{w}^{2}}=\frac{\sigma_{w}}{\sqrt{2\pi}}.\end{split}$ (81)
The first notable difference in the case of ReLU activation is that $\left\langle{U^{(1;1)}}\right\rangle\neq 0$, thus indicating the emergence of IGB. Specifically, from Eq. (74), it follows:
$f_{\mu_{c}}\left({m}\right)=\mathcal{N}\left(m;0,\sigma_{w}^{2}\frac{\sigma_{w}^{2}}{2\pi}\right)=\mathcal{N}\left(m;0,\frac{2}{\pi}\right),$ (82)
@@
In this scenario, unlike the linear activation function case discussed at the end of App. D.2, we do not need to consider the finite size effects of the dataset. While for a linear activation function $\lim_{D\rightarrow\infty}\left\langle{U^{(1;1)}}\right\rangle=0$, for ReLU $\lim_{D\rightarrow\infty}\left\langle{U^{(1;1)}}\right\rangle\neq 0$ (as shown in (81)). Therefore, the finite size corrections of $\mathcal{O}\left(\frac{1}{D}\right)$ are negligible when $D\left\langle{U^{(1;1)}}\right\rangle\gg 1$.
Figure D.1: Comparison of the distribution $f_{\mu_{c}}\left({x}\right)$ between the linear activation function (upper plot) and ReLU (bottom plot) across different dataset sizes ($D$). Both plots also include the theoretical distributions derived in App. D.2. Observe the distinct behaviors in the two cases: for the linear activation function, the distribution narrows as $D$ increases, while for the ReLU case, it remains stable. This stability in the ReLU case aligns with expectations from Eq. (81) and Eq. (77). These simulations employed the GB dataset and the SHLP model. For more details, refer to App. I.4.
In Fig. D.1, the distributions $f_{\mu_{c}}\left({m}\right)$ along with their respective theoretical curves, i.e., Eq. (78) and Eq. (82), are compared for the two types of activation functions. Consistent with Eq. (78) and Eq. (82), we observe that in the linear case, the distribution narrows around 0 as $D$ increases, whereas in the ReLU case, it remains stable.
@@
With max pooling, we group the set $\{H^{(1)}_{k}\}$ into subgroups and select the maximum value from each, thereby reducing the dimensionality of the layer. Incorporating this into the analysis in App. D.2, we have:
$U^{(1;k)}_{l}=U^{(k)}\left({\left\{H\left({Z^{(1)}_{j}}\right)\right\}_{j\in S_{l}^{k}}}\right)=\max_{j\in S_{l}^{k}}\{\max\{0,H^{(1)}_{j}\}\}$, (83)
where $S_{l}^{k}$ indicates the $l$-th subgroup of $k$ nodes.
To proceed with the computation from the previous section, we need to derive $f_{U^{(1;k)}_{l}}^{(\chi)}\left({u{}{}}\right)$. Generally, if $Y\equiv\max\{X_{1},\dots,X_{m}\}$ with $\{X_{i}\}_{i=1}^{m}$ being i.i.d., we have:
@@
$f_{Y}\left({y}\right)=m\,f_{X}\left({y}\right)F_{X}(y)^{m-1}.$ (85)
For our case, we substitute $f_{X}\left({x}\right)$ with the distribution used in App. D.2, i.e.,
$f_{X}\left({x}\right)\equiv f_{H^{(1)}_{j}}^{(\chi)}\left({x}\right)=\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right).$ (86)
@@
$\begin{split}&F_{X}(y)\equiv\mathbb{P}\left(H^{(1)}_{j}\leq y\mid\mathcal{W}\right)=\int_{-\infty}^{y}\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right)dx=\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{y}{\sqrt{2\sigma_{w}^{2}}}\right)\Theta(y).\end{split}$ (87)
Applying the Fisher–Tippett–Gnedenko theorem, we can find an asymptotic result for $f_{U^{(1;k)}_{l}}^{(\chi)}\left({u{}{}}\right)$ as $k\rightarrow\infty$. For general $k$ values, combining all components yields:
$\begin{split}&\left\langle{U^{(1;k)}}\right\rangle=\left\langle{U^{(1;k)}_{l}}\right\rangle=\int_{\mathbb{R}}u{}{}\,f_{U^{(1;k)}_{l}}^{(\chi)}\left({u{}{}}\right)du{}{}=\int_{0}^{\infty}m\,u{}{}\,\mathcal{N}\left(u{}{};0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{u{}{}}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{m-1}du{}{}.\end{split}$ (88)
This can be approximated by an asymptotic expansion or evaluated numerically. In the case of linear activation with max pooling, a simple Gaussian distribution is substituted for $f_{X}\left({x}\right)$.
@@
$c_{k}(\sigma_{w})\equiv\int_{0}^{\infty}k\,u{}{}\,\mathcal{N}\left(u{}{};0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{x}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{k-1}du{}{}.$ (89)
@@
$f_{\mu_{c}}\left({m}\right)=\mathcal{N}\left(m;0,\sigma_{w}^{2}c_{k}(\sigma_{w})^{2}\right).$ (90)
D.3 Derivation of $\boldsymbol{f_{O_{c}}^{(\chi)}\left({o}\right)}$ for some emblematic cases
As discussed, $\mu_{c}$ is a r.v. that varies with the set of weights $\mathcal{W}$. In App. D.2, we derived its distribution. In this section, we focus on the distribution of the output, given a specific configuration of the weight set, $f_{O_{c}}^{(\chi)}\left({o}\right)$ , for the same cases analyzed in App. D.2.
@@ D.3 Derivation of $\boldsymbol{f_{O_{c}}^{(\chi)}\left({o}\right)}$ for some emblematic cases
    $\lim_{N_{1}\rightarrow\infty}f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;0,\sigma_{w}^{4}\right).$ (91)
@@
    $\left\langle{H^{(1)}_{i}}\right\rangle$ $=\frac{\sigma_{w}}{\sqrt{2\pi}},$ (92)
    $\begin{split}\textrm{Var}_{\chi}\left({H^{(1)}_{i}}\right)&=\int_{\mathbb{R}^{+}}h{}{}^{2}f_{H^{(1)}_{i}}^{(\chi)}\left({h{}{}}\right)dh{}{}-\left\langle{H^{(1)}_{i}}\right\rangle^{2}=\frac{\sigma_{w}^{2}(\pi-1)}{2\pi},\end{split}$ (93)
    where the symmetry of $f_{Z^{(1)}_{k}}^{(\chi)}\left({z{}{}}\right)$ was used. Substituting into Eq. (72) and using Eq. (93), we obtain
@@
  • ReLU + max pool: This case is conceptually similar to the ReLU case; the main difference is that we do not have an analytical expression for $\left\langle{U^{(1;k)}_{i}}\right\rangle$ and $\textrm{Var}_{\chi}\left({U^{(1;k)}_{i}}\right)$. We can numerically compute these cumulants for a fixed $k$, with their expressions given by
    $\begin{split}&\left\langle{U^{(1;k)}}\right\rangle=\int_{0}^{\infty}ku{}{}\mathcal{N}\left(u{}{};0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{u{}{}}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{k-1}du{}{},\end{split}$ (95)
    $\begin{split}&\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)=\left(\int_{0}^{\infty}ku{}{}^{2}\mathcal{N}\left(u{}{};0,\sigma_{w}^{2}\right)\left(\frac{1}{2}+\frac{1}{2}\operatorname{erf}\left(\frac{u{}{}}{\sqrt{2\sigma_{w}^{2}}}\right)\right)^{k-1}dx\right)-\left\langle{U^{(1;k)}}\right\rangle^{2}.\end{split}$ (96)
@@
    $\begin{split}&\lim_{N_{1}\rightarrow\infty}f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;\mu_{c},\sigma_{w}^{2}\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)\right),\end{split}$ (97)
@@ D.4 Non identically distributed classes
$\textrm{Var}_{\chi}(O^{(c^{\prime},c)})=\frac{\sigma_{w}^{4}\sigma_{\xi^{(c)}}^{2}(\pi-1)}{2\pi}$
@@ Appendix E Single hidden layer perceptron
In this section, we utilize the results derived in App. D to deduce our primary goal, $f_{G_{0}}\left({g_{0}}\right)$. We specifically focus on a single hidden layer perceptron; in App. G.3 we will extend the results to deep architectures.
E.1 Derivation of $\boldsymbol{f_{G_{0}}\left({g_{0}}\right)}$
Consider a binary classification problem using a single hidden layer perceptron. We will use the notation outlined in Fig. A.1. The nodes of the hidden layer, after processing through the activation function and the pooling layer, are connected to two output nodes $\left\{O_{c}\right\}_{c=0}^{1}$ via two random vectors $\left\{w^{(1)}_{c\;\boldsymbol{\cdot}}\right\}_{c=0}^{1}$ . Each output node $O_{c}\left(\boldsymbol{\xi};\mathcal{W}^{0},w^{(1)}_{c\;\boldsymbol{\cdot}}\right)$ follows the distributions derived in App. D.3.
Note that, given a set $\mathcal{W}$, different output nodes depend on distinct subsets of $\mathcal{W}$, as the final matrix of weights is not shared among all nodes. The notation $O_{c}\left(\boldsymbol{\xi};\mathcal{W}^{0},w^{(1)}_{c\;\boldsymbol{\cdot}}\right)$ highlights this aspect, underscoring why the output nodes have different distributions. However, for brevity, we will denote a general dependency on the entire set $\mathcal{W}$, i.e., we will use the notation $O_{c}\left(\boldsymbol{\xi};\mathcal{W}\right)$ .
To derive the distribution $f_{G_{0}}\left({g_{0}}\right)$ , we proceed as follows:
In a single experiment, the set $\left\{w^{(1)}_{c\;\boldsymbol{\cdot}}\right\}_{c=0}^{1}$ is fixed, allowing us to compute $f_{O_{c}}^{(\chi)}\left({o}\right)$ . Recalling that $G_{0}$ represents the fraction of the data-points guessed as members of class 0, according to the Law of Large Numbers, this fraction will converge to the probability
$\lim_{D\rightarrow\infty}G_{0}\left(\mathcal{W}\right)=\mathbb{P}\left(O_{0}>O_{1}\mid\mathcal{W}\right)=\mathbb{P}\left(O_{0}-O_{1}>0\mid\mathcal{W}\right)=\int_{0}^{\infty}f_{\Delta_{O}}^{(\chi)}\left({\delta_{o}}\right)\;d\delta_{o},$ (99)
where $G_{0}\left(\mathcal{W}\right)$ indicates the fraction $G_{0}$ obtained for the specific fixed configuration $\mathcal{W}$.
$G_{0}$ is a quantity clearly dependent on the set of random variables $\mathcal{W}$. Consequently, it is a random variable itself, varying with the network’s set of weights. Defining $\Delta_{O}=O_{0}-O_{1}$ , we can express:
$f_{G_{0}}\left({g_{0}}\right)=\int_{\Omega}f_{\mathcal{W}}\left({\tilde{\mathcal{W}}}\right)\delta\left(\mathbb{P}\left(\Delta_{O}>0\mid\tilde{\mathcal{W}}\right)-g_{0}\right)d\tilde{\mathcal{W}}$ (100)